1. Calculate pH and % ionization of 0.25 M Naco (Ka (HCO2H) 1.8 x 10) 2- Calculate the pH of: a) 0.35 M HNO b) 0.15 M Sr(O)2 c) 0.08 M Ba(CIO4)2 3- Calculate the pH of 0.375 L buffer solution made of a 0.18 M Acetic acid HC2H:02(Ka 1.8x 10) and a 0.134 M Potassium acetate KC2Hs02 (do not use more than 3 digits beyond the decimal point) a) before adding anything b) after adding 0.010 mol Ba(OH)2 c) after adding 0.050 mol HCI0 following salt solutions a) 0.1 M KI(aq) b) 0.1 M NH NO (aq) c) 0.1 M Zn(CIO4)2(aq) Consider the titration of 50.0 mL of 0.200 M hypochlorous acid HCIO (Ka 3.5 x 108) with 0.250 M NaOH 5- How many milliliters of NaOH are required to reach the equivalence point? a) b) Calculate the pH before titration c) Calculate the pH after adding 40.0 mL, of NaOH d) Calculate the pOH after adding 20.0 ml. of NaOH
Answer #1

According to HomeworkLib rules we have to answer only first one problem

HCOONa >> HCOO- + Na+

HCOO- + H2O <----> HCOOH + OH-
K = Kw/Ka = 1 x 10^-14 / 1.8 x 10^-4 = x^2/ 0.25-x

x = [OH-]= 3.73 x 10^-6 M
pOH = 5.43
pH =14- p OH = 8.57

OR

pH=7+0.5 [pKa+log(C) ]

where C is the concentration of the salt.

so pH=7+ 0.5*3.72 +0.5 log(0.25)

= 7+ 1.86 -0.3010

=8.559

Percentage ionization = [OH-] / B *100

= 3.73 x 10^-6/0.25 *100

= 0.001492%

=0.0015 %